OpenStudy (anonymous):

Solve for x, rounded to the nearest thousandth. 15=5^(x+3)-7

OpenStudy (anonymous):

@satellite73, would you mind explaining how to solve this?

OpenStudy (anonymous):

15+7=5^(x+3) 22=5^(x+3) have u learnt the log/ln rules yet?